### v2

# Negativity Percolation in Continuous-Variable Quantum Networks

Abstract
Abstract
Quantum networks (QNs) have been predominantly driven by discrete‑variable (DV) architectures. Yet, optical platforms naturally generate Gaussian states—the common states of continuous-variable (CV) systems, making CV-based QNs an attractive route toward scalable, chip‑integrated quantum computation and communication. To bridge the gap between well‑studied DV entanglement percolation theories and their CV counterpart, we introduce a Gaussian-to-Gaussian entanglement distribution scheme that deterministically transports two‑mode squeezed vacuum states across large CV networks. Analysis of the scheme’s collective behavior using statistical-physics methods reveals a new form of entanglement percolation—negativity percolation theory (NegPT)—characterized by a bounded entanglement measure called the ratio negativity. We discover that NegPT exhibits a mixed-order phase transition, marked simultaneously by both an abrupt change in global entanglement and a long-range correlation between nodes. This distinctive behavior places CV-based QNs in a new universality class, fundamentally distinct from DV systems. Additionally, the abruptness of this transition introduces a critical vulnerability of CV-based QNs: conventional feedback mechanism becomes inherently unstable near the threshold, highlighting practical implications for stabilizing large-scale CV-based QNs. Our results unify statistical models for CV-based entanglement distribution and uncover previously unexplored critical phenomena unique to CV systems, providing valuable insights and guidelines essential for developing robust, feedback-stabilized QNs.
Introduction
Table 1: Comparison of percolation theories (PTs) for entanglement distribution on discrete-variable (DV) and continuous-variable (CV) quantum networks (QNs). The PTs are based on distinct entanglement measures: concurrence [49, 45] for DV qudit systems, and ratio negativity [97] for CV infinite-dimensional squeezed states.
  Entanglement distribution scheme PT / entanglement measure Phase transition type
DV Deterministic entanglement transmission (DET) ConPT / Concurrence $c$ Continuous (second order)
CV Gaussian-to-Gaussian (G-G) DET NegPT / Ratio negativity $\chi$ Discontinuous (mixed order)
Quantum information technologies [74] fundamentally rely on our ability to distribute entanglement [10, 23], the process of establishing quantum correlations between distant nodes across network architectures, from chips to the internet. This capability empowers applications ranging from quantum computation [24, 50] to secure quantum communication [11, 18, 79, 72, 91, 93, 99]. The collective behavior of entanglement distribution across large-scale systems is often described by the quantum network (QN) model [1]. Within this framework, a QN is represented as a graph where nodes represent spatially separated parties, and links denote bipartite entangled states shared between adjacent nodes. The distribution process itself is executed via protocols based on local operations and classical communication (LOCC), such as entanglement swapping [100, 15] and entanglement concentration [9, 70], along with more sophisticated schemes [6, 19, 59, 75], applied across the network’s links to establish long-range correlations.
Investigating how LOCC schemes distribute entanglement across a QN has uncovered fundamental correspondences to percolation theories. The ability to establish long-range entanglement in QN often mirrors the emergence of large-scale connectivity described by percolation theories—the statistical physics framework governing phase transitions on graphs [54, 21, 25, 13, 4, 12, 60]. This conceptual bridge between entanglement distribution and percolation was first built in 2007, linking a probabilistic entanglement distribution scheme [1] to classical bond percolation [35]. Later, the development of a more effective deterministic entanglement transmission (DET) scheme [65] (where LOCC steps succeed deterministically, eliminating classical randomness unlike earlier probabilistic schemes) led to its mapping onto an essentially distinct variant, concurrence percolation theory (ConPT) [66, 63]. This interplay continues with more recent mappings identified between quantum memories and continuum percolation [68], and between entanglement routing and path percolation [55, 67], underscoring how incorporating quantum resources like memories or repeaters may also profoundly alter the underlying percolation landscape.
These established mappings are, however, exclusively confined to discrete-variable (DV) systems (e.g., qubits). They all overlook an alternative architecture particularly prominent in optical settings: the continuous-variable (CV) system, leveraging the continuous degrees of freedom of light fields (amplitude and phase) [17, 84, 85, 81]. In contrast to DV encoding, which is inherently reliant on random single-photon sources, the CV encoding is more advantageous with its unconditional, consistent generation of entanglement via nonlinear optical interactions, bypassing the hurdles of unpredictability in DV sources. This nonrandom nature unlocks considerable potential for the scalability of quantum optical applications [20, 29, 41, 30], a potential now being realized with the advent of CV photonic chips [53].
Even with their clear advantages and growing experimental relevance, the collective characteristics of CV systems—particularly as modeled by QN—remain unexplored. Crucial questions persist: How to distribute entanglement across a CV-encoded QN? Does a unifying correspondence to percolation theory also apply? If so, does the continuous nature of CV lead to fundamentally distinct network physics compared to DV systems?
To address these conceptual gaps, here we introduce a Gaussian-to-Gaussian (G-G) DET scheme, wherein both the inputs and outputs are two-mode squeezed vacuum states (TMSVSs), defined as $|\psi^{r}\rangle=\sqrt{1-\tanh^{2}r}\sum\nolimits_{n=0}^{\infty}\tanh^{n}r|nn\rangle$, which are a class of entangled CV Gaussian states characterized by the squeezing parameter $r\in[0,\infty)$ [17, 84, 85]. Akin to DV-based DET [65], the G-G DET employs deterministic entanglement swapping [82] and concentration [51] adapted for TMSVSs, offering an avenue for entanglement distribution in optical systems.
We find that the G-G DET scheme corresponds to a new variant of percolation: the negativity percolation theory (NegPT). The theory is distinct from its DV counterparts: we define it neither on probability [35] nor concurrence [90], but on a bounded entanglement measure $\chi_{\mathcal{N}}\in[0,1]$ termed ratio negativity [97] that simplifies to $\chi\equiv\tanh r$ for TMSVS $|\psi^{r}\rangle$. This presents a broader spectrum of correspondences between entanglement distribution schemes and percolation theories (Table 1).
Further, we show that the NegPT exhibits a mixed-order phase transition [71], characterized by both a discontinuous change at a critical threshold $\chi_{\text{th}}$ and a divergent correlation length $l^{*}\sim\left|\chi-\chi_{\text{th}}\right|^{-z\nu}$. While mixed-order phase transitions have been widely observed in spin systems [71], interdependent networks [98], superconducting devices [14], overload systems [77], $k$-core percolation [92], and colloidal crystals [3], their observation in QNs is unprecedented. Our discovery extends the reach of mixed-order phase transitions beyond traditional statistical physics systems, into the rapidly evolving field of quantum information for the first time. Indeed, our finding of the NegPT, with a unique thermal critical exponent $z\nu\approx 1/2$ in Bethe lattices, clearly highlights a universality class that is distinct from both classical and concurrence percolation (where $z\nu\approx 1$) [66, 63]. We also notice that the onset of mixed-order transitions may persist for generalized CV operations and even non-Gaussian states, suggesting that the NegPT may underlie the critical phenomena for generic CV-based QNs.
Moreover, our theoretical analysis reveals a potential practical vulnerability within large-scale CV-based QNs: We show that the sharp transition at $\chi_{\text{th}}$ challenges stabilization via standard feedback, risking unstable “on/off” oscillations near the threshold—in stark contrast with continuous transitions. Given the prevalence of quantum feedback control [88, 31, 61, 62, 47, 52, 26, 32, 95, 80, 96, 64] in quantum optical implementations, this instability acts as a crucial alert: scaling CV systems requires careful feedback strategies to maintain robust operation near the “brink of failure.” We initiate the characterization of this behavior and its network implications.
Results
.1 Gaussian QN
To begin with, we consider a Gaussian QN built by bipartite, pure Gaussian states $|\Psi\rangle_{AB}$ involving multiple modes (say, $K_{A}$ modes + $K_{B}$ modes) in general. A key insight for Gaussian states is that $|\Psi\rangle_{AB}$ can be mapped into a tensor product of $K{\color[rgb]{0,0,0}\leq}\min{(K_{A},K_{B})}$ TMSVSs (along with trivial vacuum states) via local Gaussian unitaries [16]: $|\Psi\rangle_{AB}=\bigotimes_{k=1}^{K}|\psi^{r_{k}}\rangle\otimes|00\dots\rangle$, where each $r_{k}$ corresponds to the squeezing parameter of the $k$-th TMSVS. This mode decomposition [16] simplifies our QN model to primarily focus on TMSVSs, such that each TMSVS represents a QN link between two nodes. A bipartite multi-mode pure Gaussian state can be treated as multi-links that connect the same pair of nodes multiple times, which are allowed in the formation of the QN.
Figure 1: Gaussian-to-Gaussian deterministic entanglement transmission (G-G DET) scheme. Applicable to Gaussian quantum networks (QN), the scheme consists of two LOCC protocols: (a) Entanglement swapping, facilitated by homodyne detection and displacement [82]; (b) Entanglement concentration, facilitated by non-standard optical components [51]. Both protocols are deterministic, taking two (or more) TMSVS $|\psi^{r_{i}}\rangle$ as input and a new TMSVS $|\psi^{r}\rangle$ as output. (c) The two LOCC protocols map to series and parallel rules, respectively, to construct G-G DET. (d) Consider a QN example built upon three node. The G-G DET scheme consists of two steps: First, the parallel rule converts the states $|\psi^{r_{1}}\rangle$ and $|\psi^{r_{2}}\rangle$ ($r_{1}\geq r_{2}$) into $|\psi^{r_{1,2}}\rangle$ with $\sinh r_{1,2}=\sinh r_{1}\cosh r_{2}$ between $S$ and $R$; second, the series rule transforms $|\psi^{r_{1,2}}\rangle$ and $|\psi^{r_{3}}\rangle$ to the final state $|\psi^{r}\rangle$ with the ratio negativity $\text{X}_{\text{SC}}=\tanh r_{1,2}\tanh r_{3}$ between $S$ and $T$.
.2 A DET scheme for series-parallel Gaussian QNs
Our G-G DET scheme, following and extending the original DET scheme proposed for DV systems [65], consists of entanglement swapping and concentration operations applied to Gaussian states. The two operations, taking TMSVSs as the input and a new TMSVS as the output, are inherently deterministic in the CV domain. These operations can be intuitively understood as “series and parallel rules,” as commonly used in studying the effective resistance in circuit networks [65]. Hence, our G-G DET scheme can be applied to any network with a series-parallel topology [33]. The corresponding series (swapping) and parallel (concentration) rules are introduced as follows:
First, we introduce the G-G entanglement swapping. Consider a setup involving three spatially separated parties: the Source ($S$), Relay ($R$), and Target ($T$). A TMSVS $|\psi^{r_{1}}\rangle$ is shared between $S$ and $R$, while another $|\psi^{r_{2}}\rangle$ is shared between $R$ and $T$ [Fig. 1(a)]. The two parties $S$ and $T$ do not directly share any entangled state. Nevertheless, executing homodyne measurement at $R$ and applying matching displacements at $S$ and $T$ can deterministically convert $|\psi^{r_{1}}\rangle$ and $|\psi^{r_{2}}\rangle$ into a new TMSVS, $|\psi^{r}\rangle$, directly shared between $S$ and $T$ [82]. The new squeezing parameter $r$ is (ideally) given by $\tanh{r}=\tanh{r_{1}}\tanh{r_{2}}$ [82]. This can be extended to a one-dimensional chain consisting of a series of $N$ links corresponding to TMSVSs $|\psi^{r_{1}}\rangle,|\psi^{r_{2}}\rangle,\cdots,|\psi^{r_{N}}\rangle$ (see Methods), respectively, leading to the series rule [Fig. 1(c)] for $|\psi^{r}\rangle$ being the final created TMSVS.
Note that the series rule is independent of the order in which the links are transformed. This commutativity, as we will see, does not apply to the parallel rule below.
Table 2: Connectivity rules for the NegPT on different topologies. Each link has identical ratio negativity $\chi\equiv\tanh r$.
  Exact rules         Approximate rules (by the star-mesh transform [66, 63])
Topology: Series Parallel Wheatstone bridgea Kelvin bridgea (higher-order topologies)
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] $\dots$
$\mathrm{X}_{\text{SC}}$ $\chi^{N} =$${\chi}\left[\chi^{2}+\left(1-\chi^{2}\right)^{K}\right]^{-1/2}$ $\approx b(\chi,y(\chi))$ $\approx b(\chi,y(\chi\delta(\chi)))$ $\dots$
  • a
    $b(\chi,y)=\frac{\chi^{2}}{\sqrt{\chi^{4}+\left(1-y^{2}\right)\left[\left(\chi^{2}y^{2}-y^{2}+1\right)^{2}-\chi^{4}\right]}}$; $y(x)$ and $\delta(x)$ are the Y-$\Delta$ and $\Delta$-Y transforms of $x$, respectively (see Methods).
We then give the G-G entanglement concentration. It is known that Gaussian states cannot be feasibly concentrated using standard optical components [34]. The only exception we know of is provided in Ref. [51], offering a conceptual construction of LOCC for concentrating entanglement from pure Gaussian states. Such LOCC must involve non-standard optical components. Assuming this, if we consider a tensor product of two TMSVSs shared by two parties $S$ and $T$, $|\psi^{r_{1}}\rangle\otimes|\psi^{r_{2}}\rangle$ ($r_{1}\geq r_{2}$), this state can be deterministically converted to a single TMSVS and a vacuum state, $|\psi^{r}\rangle\otimes|00\rangle$ by LOCC [Fig. 1(b)] [51], where $\sinh{r}=\sinh{r_{1}}\cosh{r_{2}}$ [51]. The result also scales to $K$ parallel links, $|\Psi\rangle_{{ST}}=\mathop{\bigotimes}\nolimits_{k=1}^{K}|\psi^{r_{k}}\rangle$ (see Methods). By iteratively applying the Gaussian concentration K1K-1 times, the multi-mode state $|\Psi\rangle_{{ST}}$ is concentrated into a single TMSVS, leading to the parallel rule [Fig. 1(c)].
The parallel rule is not commutative due to the presence of a $\sinh r_{1}$ term before other $\cosh r_{k}$ terms. We find that the optimal entanglement of $|\psi^{r}\rangle$ is achieved by placing the largest squeezing parameter $r_{1}$ into $\sinh r_{1}$ (see Methods).
.3 Negativity percolation theory (NegPT)
The G-G DET scheme provides the foundation for a percolation-like theory of “quantum connectivity,” which we term negativity percolation. We consider a homogeneous QN where each link represents a TMSVS with an identical entanglement weight $\chi$. This framework is analogous to classical bond percolation on a network where each link is active with an identical probability $p$ [8].
In classical percolation theory, the connectivity between two arbitrary sets of nodes, $S$ and $T$, can be quantified by the “sponge-crossing” probability, $P_{\text{SC}}$ [66, 63]. This quantity is the combination of probabilities over all paths connecting the two sets. Drawing a parallel, we introduce the “sponge-crossing” ratio negativity, $\mathrm{X}_{\text{SC}}$, which represents an analogous combination of all paths but is framed in terms of the ratio negativity. Specifically, consider a Gaussian QN with endpoints $S$ and $T$ that initially share no entanglement. By implementing the G-G DET scheme across the QN, $S$ and $T$ ultimately share a TMSVS $|\psi^{r}\rangle$, with entanglement established between them. The final entanglement is quantified by the ratio negativity of $|\psi^{r}\rangle$, which we refer to as the sponge-crossing ratio negativity $\mathrm{X}_{\text{SC}}$.
For both classical and negativity percolation, this aggregation of all path contributions is calculated as follows:
(1) When the network topology between $S$ and $T$ is series-parallel, $\mathrm{X}_{\text{SC}}$ can be decomposed into the iteration of the exact series and parallel rules [e.g., Fig. 1(d)]. Additionally, in a series network of $N$ TMSVSs with identical ratio negativity $\chi$, $\mathrm{X}_{\text{SC}}=\chi^{N}$ has the same form as the probability measure $p$ (or concurrence $c$) in classical (or concurrence) percolation [66, 63]. This similarity prompts us to refer to the theory as the NegPT, emphasizing on the special role that ratio negativity $\chi$ plays that is analogous to $p$ (or $c$).
(2) When the network topology between $S$ and $T$ is not series-parallel, we do not know the exact transmission rules (beyond the scope of G-G DET). However, we can approximate these higher-order rules using transforms based only on the series/parallel rules (see Methods).
Table 2 shows examples on how to calculate $\mathrm{X}_{\text{SC}}$ (exactly or approximately) for various network topologies. When the number of QN nodes goes to infinity, similar to classical percolation, we assume that a nontrivial critical threshold exists, i.e., a minimum entanglement $\chi_{\text{th}}\in[0,1]$ per link that makes $\mathrm{X}_{\text{SC}}$ exceed zero.
Figure 2: Bethe lattice. (a) A Bethe lattice of degree $k$ (i.e., each node is incident to $k$ links) and network depth $l$ (the path length from the yellow node to the red nodes). (b) The sponge-crossing ratio negativity $\mathrm{X}_{\text{SC}}$ between $S$ and $T$ for various $k$ (right panel), satisfying the power law $\mathrm{X}_{\text{SC}}-\mathrm{X}_{\text{SC}}^{+}\sim|\chi-\chi_{\text{th}}|^{0.47(5)}$ as $\chi\to\chi_{\text{th}}^{+}$ (left panel). The numerical value $0.47\pm 0.05$ is derived by a linear least-squares fit to the sixteen data points. (c) When $\chi\to\chi_{\text{th}}^{-}$, $\mathrm{X}_{\text{SC}}$ exhibits a plateau behavior until the network depth $l$ exceeds the correlation length $l^{*}$ (defined as the depth $l$ at which $\mathrm{X}_{\text{SC}}=0.5$), after which $\mathrm{X}_{\text{SC}}$ abruptly drops to zero. (d) Near the critical threshold, we observe $l^{*}\sim\left|\chi-\chi_{\text{th}}\right|^{-0.508(9)}$, indicating $z\nu\approx 1/2$.
.4 Mixed-order phase transition
We start by focusing on the Bethe lattice [8], an infinite tree-like series-parallel network where every node has an identical degree $k>2$ [Fig. 2(a)]. The root (yellow) corresponds to the set $S$, and all the nodes (red) situated at the outermost boundary are collectively regarded as the set $T$. We define the (finite) network depth $l$ as the number of links in the shortest path between $S$ and $T$ (see Supplemental Information (SI), Sec. I A).
In the $l\to\infty$ limit, the Bethe lattice manifests self-similarity: the dark and light orange shaded regions are identical, as are the dark and light green shaded regions. This enables us to derive the exact self-consistent renormalization group [22, 76] equation (SI, Sec. I B). We observe a distinctive phenomenon: as $\chi$ increases from zero, $\mathrm{X}_{\text{SC}}$ initially remains zero until $\chi=\chi_{\text{th}}$, at which point $\mathrm{X}_{\text{SC}}$ exhibits an abrupt, discontinuous jump to a positive value $\mathrm{X}_{\text{SC}}^{+}$ before continuing to increase [Fig. 2(b)]. For general $k$, we find
$\chi_{\text{th}}=\sqrt{1-(k-1)^{-\frac{k-1}{k-2}}\left(k-2\right)}.$ (1)
Moreover, at the critical threshold, $\mathrm{X}_{\text{SC}}$ jumps to
$\mathrm{X}_{\text{SC}}^{+}=\sqrt{\frac{1-T(k)}{T^{k}(k)-T(k)+1}},$ (2)
where $T(k)=\left(k-1\right)^{-1/{\left(k-2\right)}}$ (SI, Sec. I C). This is in stark contrast with classical or concurrence percolation, where the phase transition is shown to be continuous [66, 63]. Another difference is that in concurrence percolation, a nontrivial saturation point exists, above which the sponge-crossing entanglement becomes unity (even when the entanglement per link is below one) [66, 63]. Yet, there is no saturation point in the NegPT. Instead, we find $1-\mathrm{X}_{\text{SC}}\sim(1-\chi)^{k}$ when $\chi\to 1^{-}$, and $\mathrm{X}_{\text{SC}}$ becomes one if and only if $\chi=1$ (SI, Sec. I D).
The discontinuity around $\chi_{\text{th}}$ prompts us to investigate its critical nature, showing that the phase transition is, indeed, mixed order (SI, Sec. I E). Below the critical threshold ($\chi<\chi_{\text{th}}$), we find that $\mathrm{X}_{\text{SC}}$ initially remains constant as a function of $l$, forming a plateau near $\mathrm{X}_{\text{SC}}\approx\mathrm{X}_{\text{SC}}^{+}$ [Fig. 2(c)]. This plateau persists until $l$ reaches a finite value $l^{*}$, after which $\mathrm{X}_{\text{SC}}$ abruptly drops to zero. The same plateau is also observed in interdependent physical networks, signaling a dynamic characteristic of mixed-order phase transitions [98].
By analogy with classical percolation [8], this value $l^{*}$ represents the characteristic length, signifying the typical distance within which two nodes remain connected. The log-log fitting [Fig. 2(d)] further indicates that $l^{*}\sim\left|\chi-\chi_{\text{th}}\right|^{-1/2}$, suggesting that $l^{*}$ approaches infinity as $\chi\to\chi_{\text{th}}^{-}$, with a power law exponent $1/2$ that is equal to the chemical-space (“time”) thermal exponent, $z\nu$, as determined by finite-size scaling analysis [44]. Notably, $z\nu\approx 1/2$ differs from the classical and concurrence percolation theories, where $z\nu\approx 1$ for both [66, 63], indicating that the underlying dynamics of the NegPT is fundamentally unique and belongs to a different universality class.
To identify the other critical exponent, $\beta$, note that akin to concurrence percolation, the series rule of the NegPT also has a “gauge redundancy” under the change of variable $\chi\to\chi^{x}$ [66, 63]. However, this redundancy no longer leads to arbitrariness in $\beta$. Indeed, for arbitrary $x$, we always have $\left(\mathrm{X}_{\text{SC}}\right)^{x}-\left(\mathrm{X}_{\text{SC}}^{+}\right)^{x}\simeq x\left(\mathrm{X}_{\text{SC}}^{+}\right)^{x-1}\left(\mathrm{X}_{\text{SC}}-\mathrm{X}_{\text{SC}}^{+}\right)\sim\left(\chi-\chi_{\text{th}}\right)^{1/2}$ (SI, Sec. I E), indicating a fixed critical exponent $\beta=1/2$ [Fig. 2(b)] independent of the gauge redundancy $x$.
Mixed-order phase transitions also occur in classical interdependent percolation [42, 48], prompting a comparison to NegPT. Consider classical percolation in $M$ fully interdependent multilayer systems [56, 69] of the Bethe lattice topology, and let $P_{\text{SC}}^{(1)}$ represent the sponge-crossing probability across a single branch [shaded orange in Fig. 2(a)]. Per the excess-degree generating function interpretation [86], full interdependency requires that all branches in the $M$ layers must be simultaneously percolating in order to enable the percolation of the next-level branch [42, 48]. This consideration leads to a self-consistent equation (SI, Sec. I F) that also features a mixed-order jump from 0 to $P_{\text{SC}}^{(1),+}$ at the critical threshold $p=p_{\text{th}}$. As $k\to\infty$, one has $P_{\text{SC}}^{(1),+}\to 0$, which leads to the expansion (SI, Sec. I G)
$P_{\text{SC}}^{(1),+}=p_{\text{th}}\left[1-M\left(1-P_{\text{SC}}^{(1),+}\right)^{k-1}+\dots\right]$ (3)
near the critical threshold. Analogously, for the NegPT, we find that the jump of the sponge-crossing ratio negativity across a single branch, $\mathrm{X}_{\text{SC}}^{(1),+}$, satisfies (SI, Sec. I G)
$\left(\mathrm{X}_{\text{SC}}^{(1),+}\right)^{2}=\chi_{\text{th}}^{2}\left[1-\frac{1}{\left(\mathrm{X}_{\text{SC}}^{(1),+}\right)^{2}}\left(1-\left(\mathrm{X}_{\text{SC}}^{(1),+}\right)^{2}\right)^{k-1}+\dots\right]$ (4)
for $k>2$. Comparing Eqs. (3) and (4), we see that as $k\to\infty$, the NegPT mirrors classical percolation in an $(\mathrm{X}_{\text{SC}}^{(1),+})^{-2}$-layer interdependent Bethe lattice under the change of variable $\chi^{2}\to p$. In general, however, the two models are driven by fundamentally different mechanisms. While classical interdependent percolation requires an additional degree of freedom (i.e., the number of layers $M$) to induce a mixed-order transition, the NegPT does not require any such extra freedom.
To verify that the discontinuous nature of NegPT’s phase transition extends beyond the Bethe lattice, we also performed numerical simulations of entanglement percolation on a square lattice using the star-mesh transform [66]. As shown in Fig. 3(a), the curve drops abruptly already for finite side length $L$, suggesting a (weakly) discontinuous transition [40, 67] may indeed exist near the critical threshold $\chi_{\text{th}}\approx 0.715$. Moreover, we define the correlation length $\xi$ by $\chi\sim e^{-L/\xi}$ for $\chi<\chi_{\text{th}}$, measuring how fast $\chi$ approaches zero with respect to increasing length $L$ in the subcritical regime. Checking the scaling $\xi\sim|\chi-\chi_{\text{th}}|^{-\nu}$ yields an exponent $\nu=0.02(2)$ [Fig. 3(b)], a value that is distinct from $\nu\approx 1.3(3)$ for ConPT [66] and is close to zero, which is a clear signature of finite correlation length for discontinuous phase transitions [37].
.5 Generalized NegPT
To examine how the mixed-order phase transition depends on the form of the series/parallel rules in Fig. 1(c), we consider generalizing these rules. Specifically, we introduce two constant factors, $\eta_{s}$ and $\eta_{p}$, modifying the series rule to $\tanh r=\eta_{s}\tanh r_{1}\tanh r_{2}\dots$ and the parallel rule to $\sinh r=\eta_{p}\sinh r_{1}\cosh r_{2}\cosh r_{3}\dots$. These generalized rules may not be physically valid across the entire parameter range for arbitrary $\eta_{s}$ and $\eta_{p}$. However, our quantitative analysis is restricted to the region near the critical threshold, where these forms serve as a valid local approximation. On the Bethe lattice, we find that the type of the phase transition depends on $\eta_{p}$. Specifically, when $\eta_{p}<\sqrt{k-1}$, only a mixed-order phase transition can exist; when $\eta_{p}\geq\sqrt{k-1}$, only a continuous phase transition can exist (SI, Sec. II A).
What could be the possible physical origins of $\eta_{s}$ and $\eta_{p}$? First, the entanglement swapping may take a more general form [39, 43, 38], leading to $\eta_{s}\leq 1$ which reflects the generalization of homodyne measurement (SI, Sec. II B). Second, the direct-product construction of the concentration operation (see Methods) appears to rely solely on the exponential tail of TMSVS in the Fock space, $\sim\chi^{n}|nn\rangle$. Therefore, we may consider generic CV states that are non-Gaussian at low photon numbers while retaining the exponential tail at high photon numbers. This could yield $\eta_{p}\leq 1$, reflecting the arbitrariness in the low-photon coefficients (SI, Sec. II C). Interestingly, comparing this with the $\eta_{p}<\sqrt{k-1}$ criterion in the Bethe lattice example suggests that only mixed-order transitions can be observed with generic CV operations as in our settings.
Figure 3: Entanglement percolation in two-dimensional square lattices. (a) $\text{X}_{\text{SC}}$ for square lattices with different side length $L$. (b) Scaling of the correlation length $\xi$ near the critical threshold $\chi_{\text{th}}\approx 0.715$ follows $\xi\sim|\chi-\chi_{\text{th}}|^{-\nu}$, with a fitted critical exponent $\nu=0.02\pm 0.02$.
.6 Instability under feedback control
The identified abrupt transition distinguishes CV-based QNs from DV-based QNs, revealing potential instability inherent in the QN design. To illustrate this, note that in any realistic scenario, quantum states and their entanglement inevitably evolve over time due to environmental influences [87, 2, 94, 27, 89, 46, 78, 5]. As entanglement degradation is often an exponentially fast process [94, 27], we assume the ratio negativity $\chi(t)$ of each link follows an exponential decay, $\chi(t)\sim\exp{\left(-t/\tau\right)}$. This implies a decay rate proportional to the current negativity, $d\chi(t)/dt\sim-\tau^{-1}\chi(t)$, where the timescale $\tau$ characterizes the system’s noise and decoherence. Although this decay is a local process (as it occurs per link), its collective effect can rapidly compromise global functionality, driving $\mathrm{X}_{\text{SC}}\to 0$ particularly as $\chi(t)$ breaches $\chi_{\text{th}}$.
To mitigate such entanglement degradation, a standard approach may utilize measurement-based quantum feedback control [89]. This involves constantly monitoring the system’s output $\mathrm{X}_{\text{SC}}(t)$ and comparing it to a desired target value $\mathrm{X}^{\text{target}}_{\text{SC}}\gg 0$. Any difference $\epsilon(t)=\mathrm{X}^{\text{target}}_{\text{SC}}-\mathrm{X}_{\text{SC}}(t)$ triggers a feedback mechanism to actively adjust system settings, counteracting the entanglement decay and maintaining $\chi(t)$ at operational levels. We further assume that such feedback mechanism contains a time delay due to processing and transmission time across the global scale [46, 78]. Taken all assumptions together, we may approximate the full system dynamics using a first-order-plus-time-delay (FOPTD) model [5]:
$\frac{d\chi(t)}{dt}$ $=$ $-\tau^{-1}\chi(t)+u(t-T_{0}).$ (5)
Here, the first term represents decay of the ratio negativity over time and $u(t)$ signifies the feedback action applied after a delay $T_{0}$. The feedback signal $u(t)$ is determined by the observed deviation $\epsilon(t)$ (SI, Sec. III A). In a physical system, for instance, $u(t)$ might correspond to increasing the pump laser power for down-conversion sources, boosting the squeezing as well as entanglement to compensate for ongoing losses [58, 7, 28, 2].
Figure 4: Feedback stabilization of QN against entanglement decay. (a) Under the same feedback control [Eq. (5)], the DV-based QN ($k=3$ Bethe lattice) exhibits rapid stabilization; (b) whereas the CV-based QN exhibits long-term “on/off” instability, a direct result of the abrupt drop in Fig. 2(b).
Our aim is to understand how the sharp, discontinuous nature of $\mathrm{X}_{\text{SC}}$ vs. $\chi$ [Fig. 2(b)] undermines the feedback efficacy. Simulating an FOPTD model with identical $\tau$, $T_{0}$, and controller $u(t)$ on both DV- and CV-based Bethe-lattice QNs reveals starkly different behaviors. Under the concurrence percolation [66, 63], the sponge-crossing concurrence $C_{\text{SC}}(t)$ for DV-based states recovers smoothly under feedback [Fig 4(a)]. In contrast, the sponge-crossing ratio negativity $\mathrm{X}_{\text{SC}}(t)$ for CV-based states, dictated by its mixed-order transition, exhibits unstable “on/off” oscillations using the same $u(t)$ [Fig. 4(b)].
Crucially, this instability is not a result of the limitation of the FOPTD model’s capability. Indeed, under Eq. (5), $\chi(t)$ is successfully stabilized within a narrow range and only undergoes small perturbations near the critical threshold $\chi_{\text{th}}\approx 0.866$ (SI, Sec. III B). Therefore, the instability of $\mathrm{X}_{\text{SC}}$ comes directly from the sharp nature of $\mathrm{X}_{\text{SC}}$ vs. $\chi$ [Fig. 2(b)]. We also find that the instability persists across a wide range settings for $u(t)$ (SI, Sec. III C), indicating that the CV system stabilization is substantially more demanding than for DV counterparts, calling for a more careful feedback design for CV-based QNs.
Discussion
Despite the rich studies on mixed-order phase transitions [71, 98, 14, 77, 92, 3], the underlying mechanisms remain not fully understood. The traditional perspective of spinodal points [57] seems not applicable here, since our negativity percolation model—based solely on series and parallel rules—lacks a manifestation of hysteresis or metastability (which is also commonly absent in interdependent percolation [42, 48]). Consequently, the emergence of singularity may not necessarily be described by a partition function as for spinodal points [57]. This heralds future theoretical investigations into how the interplay of series and parallel rules can induce mixed-order transitions. Such studies could unveil novel phase transition landscapes potentially observable in quantum optical experiments implementing our G-G DET scheme.
For practical QN design, several unresolved questions remain. For instance, unlike in DV systems [65], our CV parallel rule (entanglement concentration from Ref. [51]) is not optimal. Understanding the feasibility and implementation limits of entanglement concentration has been a focus for CV system designs. This understanding is becoming even more urgent as we approach the deployment of larger-scale, realistic QNs.
Methods
This section presents the foundational tools for studying negativity percolation on Gaussian quantum networks. Further details are provided in the Supplementary Information.
.1 G-G entanglement swapping
We employ Gaussian-to-Gaussian entanglement swapping for pure Gaussian states as proposed in Ref. [82]. Consider the simplest case of three nodes ($S$, $R$, and $T$) sharing four modes, labeled as 1 through 4, corresponding to one mode on $S$, two modes on $R$, and one mode on $T$. When a TMSVS with squeezing parameters $r_{1}$ is shared by $S$ and $R$ and another TMSVS with squeezing parameter $r_{2}$ is shared by $R$ and $T$, the CV Bell measurement is performed on modes $2$ and $3$ by detecting the output modes of a $50:50$ beam splitter with these two modes as the input modes. The quadratures $\hat{x}_{u}=(\hat{x}_{2}-\hat{x}_{3})/\sqrt{2}$ and $\hat{p}_{v}=(\hat{p}_{2}+\hat{p}_{3})/\sqrt{2}$ are measured, where $\hat{x}_{m}$ and $\hat{p}_{m}$ denote the “position” and “momentum” operators of mode $m$ ($m=2,3$). Then, the quadratures for modes $1$ and $4$ undergo displacements, $\hat{x}_{m}^{\prime}=\hat{x}_{m}-g_{m}\hat{x}_{u}/\sqrt{2}$ and $\hat{p}_{m}^{\prime}=\hat{p}_{m}+g_{m}\hat{p}_{v}/\sqrt{2}$ ($m=1,4$). With the right choice of the gains,
$g_{1}=\frac{2\sinh 2r_{1}}{\cosh 2r_{1}+\cosh 2r_{2}},$ (6)
$g_{4}=\frac{2\sinh 2r_{2}}{\cosh 2r_{1}+\cosh 2r_{2}},$ (7)
the output is a new TMSVS with squeezing parameter $r$ that satisfies
$\tanh r=\tanh r_{1}\tanh r_{2}.$ (8)
This protocol can be extended to a one-dimensional chain, consisting of nodes $S$, $T$, and N1N-1 intermediate nodes $R_{1},\dots,R_{N-1}$ (Fig. 5), leading to the following theorem:
Theorem 1.
Given $N$ TMSVSs $|\psi^{r_{1}}\rangle$, $|\psi^{r_{2}}\rangle$, …, $|\psi^{r_{N}}\rangle$, where $|\psi^{r_{1}}\rangle$ is shared by two parties $S$ and R1, $|\psi^{r_{2}}\rangle$ is shared by R1 and R2, …, and $|\psi^{r_{N}}\rangle$ is shared by RN-1 and $T$, the TMSVSs can be deterministically converted to a single TMSVS $|\psi^{r}\rangle$ between $S$ and $T$ by LOCC, where
$\tanh{r}=\prod_{j=1}^{N}\tanh{r_{j}}.$ (9)
Notably, the final result is independent of the order of execution. This property, as we will see, does not apply to the concentration protocol.
We may also express Eq. (9) in terms of the ratio negativity. Let $\chi_{j}\in[0,1]$ be the ratio negativity of each input TMSVS $|\psi^{r_{j}}\rangle$, and let $\chi\in[0,1]$ be the ratio negativity of the final output TMSVS $|\psi^{r}\rangle$. We define the series function
$\text{seri}(\chi_{1},\chi_{2},...):=\chi=\prod^{N}_{j=1}\chi_{j},$ (10)
which corresponds to the series rule for the G-G DET.
Figure 5: Continuous-variable entanglement swapping. Given $N$ TMSVSs $|\psi^{r_{1}}\rangle,|\psi^{r_{2}}\rangle,\dots,|\psi^{r_{N}}\rangle$, the final state $|\psi^{r}\rangle$ between $S$ and $T$ derived by entanglement swapping is a TMSVS, where the squeezing parameter $r$ satisfies Eq. (9).
.2 G-G entanglement concentration
In DV systems, the process of entanglement concentration enables the deterministic conversion of two (or more) bipartite pure states into a single bipartite pure state with enhanced entanglement through LOCC [9]. We aim to do the same for TMSVSs, deterministically concentrating two (or more) TMSVSs into another one with higher entanglement (larger squeezing parameter). We begin with some preliminaries on deterministic LOCC:
We now introduce the majorization theory between infinite-dimensional vectors. Let $\vec{x}$ and $\vec{y}$ be positive (each element is a non-negative real number), infinite-dimensional vectors. $\vec{x}$ is majorized by $\vec{y}$, denoted as $\vec{x}\prec\vec{y}$, iff
$\Sigma_{j=1}^{k}x_{j}^{\downarrow}\leq\Sigma_{j=1}^{k}y_{j}^{\downarrow}$ (11)
holds for all positive integers $k$ and the equality holds for $k=\infty$. Here, $x_{j}^{\downarrow}$ ($y_{j}^{\downarrow}$) is the $j$th element of vector $\vec{x}$ ($\vec{y}$) arranged in descending order. Majorization theory implies that a bipartite pure state $|\psi\rangle$ can be deterministically transformed into another state $|\psi^{\prime}\rangle$ using LOCC iff $\vec{\lambda}\prec\vec{\lambda}^{\prime}$, where $\vec{\lambda}$ and $\vec{\lambda^{\prime}}$ are the Schmidt-value vectors of $|\psi\rangle$ and $|\psi^{\prime}\rangle$, with $\vec{\lambda}=(\lambda_{1},\lambda_{2},\cdots)^{\mathrm{T}}$ and $\vec{\lambda^{\prime}}=(\lambda_{1}^{\prime},\lambda_{2}^{\prime},\cdots)^{\mathrm{T}}$ [73] where $\sum_{j}\lambda_{j}=\sum_{j}\lambda_{j}^{\prime}=1$. For a TMSVS $|\psi^{r}\rangle$, its Schmidt-value vector $\vec{\lambda}(r)$ has entries $\lambda_{n}=\left(1-\tanh^{2}r\right)\tanh^{2n}(r)$. The majorization relation $\vec{x}\prec\vec{y}$ holds iff there exists a column-stochastic matrix $\mathbf{D}$ (whose column sum =1=1 for each column, and row sum $\leq 1$ for each row), such that $\vec{x}=\mathbf{D}\vec{y}$. Therefore, to construct deterministic LOCC for two TMSVSs with Schmidt-value vectors $\vec{x}$ and $\vec{y}$, respectively, conceptually one only needs to find a column-stochastic matrix $\mathbf{D}$ that satisfies $\vec{x}=\mathbf{D}\vec{y}$.
Here, we adopt the exact construction of $\mathbf{D}$ given in Ref. [51], where $\mathbf{D}$ is constructed as a tensor product: $\mathbf{D}=\mathbf{D}^{\mathcal{L}}\otimes\mathbf{D}^{\mathcal{A}}$. The matrix $\mathbf{D}^{\mathcal{L}}$ is given by
$\mathbf{D}^{\mathcal{L}}=\begin{bmatrix}1&1-\eta&(1-\eta)^{2}&(1-\eta)^{3}&\dots\\ 0&{1\choose 1}\eta&{2\choose 1}\eta(1-\eta)&{3\choose 1}\eta(1-\eta)^{2}&\dots\\ 0&0&{2\choose 2}\eta^{2}&{3\choose 2}\eta^{2}(1-\eta)&\dots\\ 0&0&0&{3\choose 3}\eta^{3}&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{bmatrix}$. (12)
Similarly, the matrix $\mathbf{D}^{\mathcal{A}}$ is given by
$\mathbf{D}^{\mathcal{A}}=\begin{bmatrix}a&0&0&0&\dots\\ ab&a^{2}&0&0&\dots\\ ab^{2}&{2\choose 1}a^{2}b&a^{3}&0&\dots\\ ab^{3}&{3\choose 1}a^{2}b^{2}&{3\choose 2}a^{3}b&a^{4}&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{bmatrix}$, (13)
where $G>0$, $a=1/G$ and b=1ab=1-a.
Note that the forms of these matrices are, indeed, motivated by the operation of a pure-loss channel with transmissivity $\eta\in(0,1)$ and a quantum-limited amplifier with a gain $G>1$ for thermal lights, respectively. This physical analogy may provide some insights on how to construct the exact LOCC using non-standard optical components.
The introduction of these two matrices is particularly useful for TMSVS, since their tensor product $\mathbf{D}^{\mathcal{L}}\otimes\mathbf{D}^{\mathcal{A}}$ can transform the Schmidt-value vector of the tensor product of two TMSVSs, $\vec{\lambda}(r_{1}^{\prime})\otimes\vec{\lambda}(r_{2}^{\prime})$, into another vector of the same form $\vec{\lambda}(r_{1})\otimes\vec{\lambda}(r_{2})$, where $\tanh r_{1}^{\prime}=\tanh r_{1}/\sqrt{\eta+(1-\eta)\tanh^{2}r_{1}}$ and $\tanh r_{2}^{\prime}=\sqrt{1-G(1-\tanh^{2}r_{2})}$. The majorization relation $\vec{\lambda}(r_{1})\otimes\vec{\lambda}(r_{2})\prec\vec{\lambda}(r_{1}^{\prime})\otimes\vec{\lambda}(r_{2}^{\prime})$ holds if $\mathbf{D}^{\mathcal{L}}\otimes\mathbf{D}^{\mathcal{A}}$ is column-stochastic, which is true iff $\eta G\geq 1$. This gives rise to the following Lemma:
Lemma 1.
(Theorem 1 in Ref. [51]) Let $|\Psi\rangle$, $|\Psi^{\prime}\rangle$ be two $(2+2)$-mode pure Gaussian states, with decreasingly ordered squeezing vectors $\vec{r}^{\downarrow}=(r_{1},r_{2})$ and $\vec{r^{\prime}}^{\downarrow}=(r_{1}^{\prime},r_{2}^{\prime})$ such that the two components of vector $(r_{1}^{\prime}-r_{1},r_{2}^{\prime}-r_{2})$ have opposite signs. Then $|\Psi\rangle$ can be transformed into $|\Psi^{\prime}\rangle$ using deterministic non-Gaussian LOCC if
$\frac{\sinh(r_{1}+r_{2})\pm\sinh(r_{1}-r_{2})}{\sinh(r_{1}^{\prime}+r_{2}^{\prime})\pm\sinh(r_{1}^{\prime}-r_{2}^{\prime})}\geq 1,$ (14)
where $\pm$ follows the sign of $r_{1}^{\prime}-r_{1}$.
Lemma 1 provides a sufficient condition to convert two TMSVSs $|\psi^{r_{1}}\rangle$ and $|\psi^{r_{2}}\rangle$ into another one with squeezing parameter $r_{1}^{\prime}>r_{1}$ for the case of $r_{2}^{\prime}=0$. To be exact, the maximum value $r$ of $r_{1}^{\prime}$ that satisfies Eq. (14) is given by $r=\sinh^{\mbox{-}1}(\sinh{r_{1}}\cosh{r_{2}})>r_{1}$ and $r_{2}^{\prime}=0$ with $G=\eta^{-1}=(1-\tanh r_{2})^{-1}$, implying that entanglement is concentrated. Thus, consider a $(2+2)$-mode pure Gaussian state $|\psi^{r_{1}}\rangle\otimes|\psi^{r_{2}}\rangle$ with decreasingly ordered squeezing vectors $\vec{r}^{\downarrow}=(r_{1},r_{2})$. This state can be deterministically converted into a single TMSVS $|\psi^{r}\rangle$ by non-Gaussian LOCC, where
$\sinh{r}=\sinh{r_{1}}\cosh{r_{2}}.$ (15)
That is, according to Ref. [51], the bipartite state $|\psi^{r_{1}}\rangle\otimes|\psi^{r_{2}}\rangle$ can be converted into $|\psi^{r}\rangle$ ($r_{1}<r$) via deterministic LOCC when the reduced state of $|\psi^{r}\rangle$ can be transformed into the reduced state of $|\psi^{r_{1}}\rangle$ through a pure loss channel with intensity transmittance $\eta=1-\tanh^{2}r_{2}$ ($r_{2}\leq r_{1}$).
Equation (15) is the building block for scaling the concentration protocol. Let $|\Psi\rangle$ be a pure Gaussian state and, w.l.o.g., write $|\Psi\rangle=\mathop{\bigotimes}\nolimits_{k=1}^{K}|\psi^{r_{k}}\rangle$ $(r_{1}\geq r_{2}\geq\cdots\geq r_{K}>0)$ with squeezing vector $\vec{r}^{\downarrow}=\left(r_{1},r_{2},\dots,r_{K}\right)$ by Ref. [16]. By iteratively applying Eq. (15) K1K-1 times on $|\Psi\rangle$ (Fig. 6), the pure Gaussian state $|\Psi\rangle$ is deterministically concentrated into a single TMSVS. Only LOCC is used throughout the entire process, giving rise to the following theorem:
Theorem 2.
A pure Gaussian state $|\Psi\rangle_{AB}=\mathop{\bigotimes}\nolimits_{k=1}^{K}|\psi^{r_{k}}\rangle_{A_{k}B_{k}}$ with $K$-dimensional squeezing parameter vector $\vec{r}^{\downarrow}=(r_{1},\dots,r_{K})$ (where $r_{1}\geq\dots\geq r_{K}>0$) can be deterministically converted to a TMSVS $|\psi^{r}\rangle$ by non-Gaussian LOCC, where $r$ satisfies the series rule
$\sinh{r}=\sinh{r_{1}}\prod_{k=2}^{K}\cosh{r_{k}}.$ (16)
We regard $|\psi\rangle$ as the tensor $\otimes_{k=1}^{K}|\psi^{r_{k}}\rangle$ of TMSVSs. Firstly, state $|\psi^{r_{1}^{\prime}}\rangle$ with
$\sinh{r_{1}^{\prime}}=\sinh{r_{1}}\cosh{r_{2}}>\sinh{r_{1}}.$
can be generated from $|\psi^{r_{1}}\rangle\otimes|\psi^{r_{2}}\rangle$ by entanglement concentration. Next, executing the scheme again on state $|\psi^{r_{1}^{\prime}}\rangle\otimes|\psi^{r_{3}}\rangle$, we get $|\psi^{r_{2}^{\prime}}\rangle$ for
$\sinh{r_{2}^{\prime}}=\sinh{r_{1}}\prod_{k=2}^{3}\cosh{r_{k}}.$
Similarly, applying the scheme iteratively, we derive a state $|\psi^{r}\rangle$ with
$\sinh{r}=\sinh{r_{1}}\prod_{k=2}^{K}\cosh{r_{k}}.$
Figure 6: Continuous-variable entanglement concentration. Consider $K$ parallel TMSVSs written as the tensor product: $\bigotimes\nolimits_{k=1}^{K}|\psi^{r_{k}}\rangle$. Step $a_{1}$: The TMSVS $|\psi^{r_{1}^{\prime}}\rangle$ is obtained by executing the TMSVS entanglement concentration protocol on $|\psi^{r_{1}}\rangle\otimes|\psi^{r_{2}}\rangle$; Step $a_{2}$: Performing the scheme again on $|\psi^{r_{1}^{\prime}}\rangle\otimes|\psi^{r_{3}}\rangle$ yields the TMSVS $|\psi^{r_{2}^{\prime}}\rangle$; and so on. This eventually results in a single TMSVS $|\psi^{r}\rangle$, where the squeezing parameter $r$ satisfies Eq. (16).
We now illustrate the optimal concentration order. In general, depending on which two modes are first concentrated, the prefactor in Eq. (16) could be $\sinh{r_{k}}$ for any $k$ other than $\sinh{r_{1}}$. However, the optimal entanglement of $|\psi^{r}\rangle$ is obtained by first concentrating the mode of the highest entanglement (largest squeezing parameter, $r_{1}$):
Theorem 3.
The squeezing parameter $r$ of the final TMSVS $|\psi^{r}\rangle$ is maximized if and only if the TMSVS with the largest squeezing parameter ($\max\limits_{1\leq k\leq K}r_{k}$) is among the first two TMSVSs to be concentrated. The corresponding optimal $r$ is given by $\sinh r=(\tanh{\max\limits_{1\leq k\leq K}r_{k})}\left(\prod_{k=1}^{K}\cosh r_{k}\right)$.
Assume w.l.o.g. $r_{1}\geq r_{2}\geq\dots\geq r_{K}$. We first consider the simplest nontrivial case $K=3$ and then generalize to $K>3$:
Case 1 ($K=3$). If $|\psi^{r_{2}}\rangle$ and $|\psi^{r_{3}}\rangle$ are concentrated first (without $|\psi^{r_{1}}\rangle$), the final squeezing parameter $r$ satisfies
$\sinh r=\max\{\cosh r_{1}\sinh r^{\prime},\ \sinh r_{1}\cosh r^{\prime}\},$
where $r^{\prime}=\text{arcsinh}(\sinh r_{2}\cosh r_{3})$. Since $r_{1}\geq r_{k}$ for $k=2,3$, we have $\sin r_{1}\cosh r_{2}\cosh r_{3}\geq\cosh r_{1}\sinh r^{\prime}$. Furthermore, $\sin r_{1}\cosh r_{2}\cosh r_{3}\geq\sinh r_{1}\cosh r^{\prime}$ holds, due to
$\cosh r^{\prime}$ $=$ $\sqrt{1+\sinh^{2}r^{\prime}}$
$=$ $\sqrt{1+\sinh^{2}r_{2}\cosh^{2}r_{3}}$
$=$ $\sqrt{\cosh^{2}r_{2}\cosh^{2}r_{3}-(\cosh^{2}r_{3}-1)}$
$\leq$ $\cosh r_{2}\cosh r_{3}.$
Thus, $r$ would be greater if $|\psi^{r_{1}}\rangle$ was concentrated first.
Case 2 ($K>3$). Assume that this theorem holds true for $K=n\ (n\geq 3)$. Then, for $K=n+1$, $|\psi^{r_{1}}\rangle$ must either be the first to concentrate (since it is the largest among the first $n$ TMSVSs to be concentrated), or the last to concentrate. In the latter case, $|\psi^{r_{2}}\rangle$ (with the second largest $r_{2}$), which is now the largest among the first $n$ TMSVSs, must be the first to be concentrated. This gives rise to
$\sinh r=\max\{\cosh r_{1}\sinh r^{\prime\prime},\ \sinh r_{1}\cosh r^{\prime\prime}\},$
where $r^{\prime\prime}=\text{arcsinh}\left(\sinh r_{2}\prod_{k=3}^{n+1}\cosh r_{k}\right)$. Because $r_{1}\geq r_{2}$ and
$\cosh r^{\prime\prime}$ $=$ $\sqrt{1+\sinh^{2}r^{\prime\prime}}$
$=$ $\sqrt{1+\sinh^{2}r_{2}\prod_{k=3}^{n+1}\cosh^{2}r_{k}}$
$=$ $\sqrt{\prod_{k=2}^{n+1}\cosh^{2}r_{k}-(\prod_{k=3}^{n+1}\cosh^{2}r_{k}-1)}$
$\leq$ $\prod_{k=2}^{n+1}\cosh r_{k},$
we have $\sinh r_{1}\prod_{k=2}^{n+1}\cosh r_{k}\geq\cosh r_{1}\sinh r^{\prime\prime}$ and $\sinh r_{1}\prod_{k=2}^{n+1}\cosh r_{k}\geq\sinh r_{1}\cosh r^{\prime\prime}$, respectively. It is therefore optimal to concentrate $|\psi^{r_{1}}\rangle$ first. ∎
To express Eq. (16) in terms of entanglement measure, let $\chi_{k}$ and $\chi$ be the ratio negativity of input TMSVS $|\psi^{r_{k}}\rangle$ and the output TMSVS $|\psi^{r}\rangle$, respectively. We have:
$\frac{\chi^{2}}{1-\chi^{2}}=\frac{\max\limits_{1\leq k\leq K}\chi_{k}^{2}}{\prod^{K}_{k=1}\left(1-\chi_{k}^{2}\right)},$ (17)
yielding a parallel function
$\text{para}(\chi_{1},\chi_{2},...):=\chi=\frac{\max\limits_{1\leq k\leq K}\chi_{k}}{\sqrt{\max\limits_{1\leq k\leq K}\chi_{k}^{2}+\prod_{k=1}^{K}(1-\chi_{k}^{2})}}.$
.3 Negativity Percolation in Non-Series-Parallel QNs
For non-series-parallel QNs, the negativity percolation theory (NegPT) based on the ratio negativity is defined using approximate star-mesh transforms (more details are provided in Ref. [66]). A star-mesh transform [83] can establish a local equivalence of the connectivity between an $(s+1)$-node star graph and an $s$-node complete graph, based only on series and parallel rules. Note that the star-mesh transform is generally irreversible. Specifically, while a star graph with arbitrary weights can most likely be transformed into a complete graph with matching weights, the inverse transformation from a complete graph with arbitrary weights back to a star graph is usually not possible. An exception to this is when $s=3$, in which case the star-mesh transform is reversible and is known as the “Y-$\Delta$$\Delta$-Y” transforms.
In the following, we examine several non-series-parallel examples (“bridge” topologies and lattices) and explain how to calculate the sponge-crossing ratio negativity $\mathrm{X}_{\text{SC}}$ using the star-mesh transform.
Figure 7: Wheatstone bridge. (a) The original QN. (c) The QN after the “Y-$\Delta$” transform. (c) The final QN.
From Eqs. (10) and (.2), we have
$\text{seri}(\chi_{1},\chi_{2})$ $=$ $\chi_{1}\chi_{2}$
and
$\text{para}(\chi_{1},\chi_{2})$ $=$ $\frac{\max\{\chi_{1},\chi_{2}\}}{\sqrt{\max\{\chi_{1}^{2},\chi_{2}^{2}\}+(1-\chi_{1}^{2})(1-\chi_{2}^{2})}}.$
The simplest non-series-parallel network is a 4-node network: a Wheatstone bridge, as shown in Fig. 7(a). Let each link in the QN be a TMSVS with identical ratio negativity $\chi$. We apply the star-mesh transform to $R_{2}$, namely the “Y-$\Delta$” transform; then, the QN becomes a series-parallel graph shown in Fig. 7(b), with three new TMSVSs of ratio negativity $\xi$ shared between $S$ and $R_{1}$, $R_{1}$ and $T$, and $S$ and $T$, respectively. Here, $\xi$ satisfies (by the star-mesh transform):
$\mathrm{seri}(\chi,\chi)=\mathrm{para}(\xi,\xi^{2}),$ (19)
namely,
$\chi^{2}=\frac{\xi}{\sqrt{\xi^{6}-\xi^{4}+1}}.$ (20)
The solution of $\xi$ is given by $\xi=y(\chi)$, where
$y(x)=\sqrt{\frac{1+2\sqrt{1+3x^{-4}}\cos{\frac{\theta+\pi}{3}}}{3}},$ (21)
and $\theta=\arccos{\left[2^{-1}(-9x^{-4}+25)(1+3x^{-4})^{-\frac{3}{2}}\right]}$. Specifically, let $W=\chi^{4}$ and $Y=\xi^{2}$. Then, Eq. (20) is equivalent to
$Y^{3}-Y^{2}-W^{-1}Y+1=0,$ (22)
where $0\leq Y\leq W\leq 1$. The Shengjin’s formula [36] implies that the roots of this cubic polynomial are
[…]
where $\theta=\arccos\left[2^{-1}(-9W^{-1}+25)(1+3W^{-1})^{-\frac{3}{2}}\right]$. Only the third root $Y_{3}$ meets the numerical requirements $0\leq Y\leq W\leq 1$. Therefore, we have $\xi=\sqrt{Y_{3}}=y(\chi)$.
Solving Fig. 7(b), the sponge-crossing ratio negativity $\mathrm{X}_{\text{SC}}$ is the ratio negativity of the final TMSVS between node $S$ and $T$ [Fig. 7(c)], which is given by $\mathrm{X}_{\text{SC}}=b(\chi,y(\chi))$ with
$b(x,y)=\frac{x^{2}}{\sqrt{x^{4}+(1-y^{2})\left[\left(x^{2}y^{2}-y^{2}+1\right)^{2}-x^{4}\right]}}.$
Figure 8: Kelvin bridge. (a) The original QN. (b) The QN after the $\Delta$-Y” transform. (c) The QN after applying series rules to $R_{1}$, $R_{2}$ and $R_{3}$. (d) The final QN.
A Kelvin bridge, the second simplest non-series-parallel network with 5 nodes [Fig. 8(a)], can be viewed as the QN after performing the star-mesh transform on node $R$ in the QN shown in Fig. 8(b), where $\delta:=y^{-1}(\chi)$ is the inverse function of $y(\chi)$. This inverse progress from Fig. 8(a) to Fig. 8(b) is thus known as the $\Delta$-Y” transform. Applying the series rule on nodes $R_{1}$, $R_{2}$ and $R_{3}$, we obtain Fig. 8(c), which is the Wheatstone bridge QN as discussed above. Consequently, $\mathrm{X}_{\text{SC}}$ for the Kelvin bridge [Fig. 8(d)] is given by
$\mathrm{X}_{\text{SC}}=b\left(\chi,y\left(\chi\delta(\chi)\right)\right).$ (25)
Data availability
Data generated and analyzed in this study are available from the corresponding authors upon request.
Code availability
Code used to generate data in this study are available from the corresponding authors upon request.
Acknowledgements
Y.Z. (Zhao), K.H., and J.H. were supported by the National Natural Science Foundation of China under Grants No. 12271394 and No. 12071336. S.H. was supported by the Israel Ministry of Innovation, Science & Technology (grant number 01017980), the Israel Science Foundation (grant number 201/25), the Binational Israel-China Science Foundation (grant number 3132/19), and the European Union under the Horizon Europe grant OMINO (grant number 101086321).
Author contributions
All authors participated in discussing the results and reviewing the manuscript. The original concept was conceived by Y.Z. (Zhao), K.H., and J.H. Numerical computations were performed by Y.Z. (Zhao) and X.M. The theoretical analysis was conducted by Y.Z. (Zhao), X.M., and Y.Z. (Zhang). The initial draft of the manuscript was written by Y.Z. (Zhao) and X.M., and revised by Y.Z. (Zhang), J.G., S.H., K.H., and J.H. All authors have read and approved the manuscript.
Competing interests
The authors declare no competing interests.